OpenStudy (anonymous):

The difference between two numbers is 28. The greater is 8 less than twice the lesser. Find the numbers.

OpenStudy (anonymous):

I have to find two equation but I found only 1 which is x-y=28

OpenStudy (anonymous):

help me plz

jhonyy9 (jhonyy9):

so let these numbers a and b what you know again ?

jhonyy9 (jhonyy9):

that a-b=28 and ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

ok what ?

jhonyy9 (jhonyy9):

you need writing an equation with what you know again

OpenStudy (anonymous):

I atill don't get it

jhonyy9 (jhonyy9):

why ? this is easy

OpenStudy (anonymous):

is x-y=28 right ?

jhonyy9 (jhonyy9):

yes this i have wrote with a and b and you know again that a>b or x>y and x-8=2y yes ?

OpenStudy (anonymous):

thank !

jhonyy9 (jhonyy9):

can you solve it now ?

OpenStudy (anonymous):

yep

jhonyy9 (jhonyy9):

ok good luck

OpenStudy (skullpatrol):

x = 2y - 8

OpenStudy (skullpatrol):

The greater is 8 less than twice the lesser.